OpenStudy (anonymous):

-(cube root of 2x^2y^2) * 2 times the cube root of 15x^5y

OpenStudy (mertsj):

\[-\sqrt[3]{2x ^{2}y ^{2}}2\sqrt[3]{15x ^{5}y}\]

OpenStudy (mertsj):

\[-2\sqrt[3]{30x ^{6}xy ^{3}?}\]

OpenStudy (mertsj):

\[-2x ^{2}y \sqrt[3]{30x}\]